# Out-of-Support Generalisation via Weight Space Sequence Modelling

### v1

@@ Abstract
As breakthroughs in deep learning transform key industries, models are increasingly required to extrapolate on datapoints found outside the range of the training set, a challenge we coin as out-of-support (OoS) generalisation. However, neural networks frequently exhibit catastrophic failure on OoS samples, yielding unrealistic but overconfident predictions. We address this challenge by reformulating the OoS generalisation problem as a sequence modelling task in the weight space, wherein the training set is partitioned into concentric shells corresponding to discrete sequential steps. Our WeightCaster framework yields plausible, interpretable, and uncertainty-aware predictions without necessitating explicit inductive biases, all the while maintaining high computational efficiency. Emprical validation on a synthetic cosine dataset and real-world air quality sensor readings demonstrates performance competitive or superior to the state-of-the-art. By enhancing reliability beyond in-distribution scenarios, these results hold significant implications for the wider adoption of artificial intelligence in safety-critical applications.
@@ 1 Introduction
Leveraging recent advances in sequence modelling and weight space learning Schürholt et al. [2024]; Nzoyem et al. [2026], we propose WeightCaster : a framework recasting OoS generalisation tasks as forecasting problems through the partitioning of the training set into concentric shells that we call “rings”. By mapping each ring to a discrete time step, we learn sequential dynamics extrapolatable to the test set. Our contributions are threefold: (1) a computationally efficient, parametric, interpretable, and inductive bias-free framework for OoS generalisation; (2) a linearisation strategy enabling the framework to provide uncertainty estimates both in-distribution (InD) and OoS; and (3) empirical validation on synthetic sinusoidal and real-world air quality experiments revealing superior or competitive performance at low parameter count.
@@ 1.2 Related Work
Our approach is also heavily connected to the nascent area of weight space learning (WSL) Schürholt et al. [2024]. Remarking on the growing capacity of public model repositories such as CivitAI and HuggingFace Jain [2022], WSL has recently captured the attention of the deep learning community, showcasing huge breakthroughs in implicit neural representations Dupont et al. [2022], generalisation error prediction Unterthiner et al. [2020], and sequence modelling Nzoyem et al. [2026].
@@ 2 Method
Fig. 1 summarises our proposed WeightCaster using a 1D sine wave as our example. We begin by selecting an anchor point from the training dataset, and we decompose the training domain into successive hyperspherical shells (equivalent to intervals in 1D or annuli in 2D). While their geometric structure might be different depending on $D_{x}$ and the chosen distance metric, we refer to these shells as “rings” for simplicity. Each ring corresponds to a step in a sequence model, and a weight space sequence model learns to predict suitable weights for each step.
Figure 1: Illustration of the two main steps of the WeightCaster framework for sinusoidal extrapolation. (a) First, an anchor point is chosen and the input domain $\mathbb{R}^{1}$ is decomposed into $T=10$ “rings”, here clearly delineated as intervals. Within each ring, we consider a simple linear model $\hat{y}=\theta^{1}\cdot x+\theta^{2}$, where $\theta=[\theta^{1},\theta^{2}]^{T}$ contains the slope and intercept to anchor, respectively. (b) Optimal weights $\{\theta_{t}\}_{t=1}^{T_{\text{tr}}}$ for the data in each ring are subsequently computed by fitting a weight space sequence model, one ring corresponding to one time step. Suitable weights for OoS datapoints are obtained by rolling out the sequence model for time steps $t\geq T_{\text{tr}}$ (in this example, $T_{\text{tr}}=9$).
In the following paragraphs of this section, we detail the ingredients of the WeightCaster framework. We begin by outlining the input domain decomposition and weight space sequence modelling strategies. Deterministic training and inference algorithms are discussed, followed by a stochastic framework for uncertainty estimation.
@@ 2.1 Domain Decomposition
Conventional machine learning seeks to learn a single model $\hat{y}=f_{\theta}(x)$, $\forall x\in X^{\text{tr}}\triangleq\bigcup_{t=1}^{T_{\text{tr}}}X_{t}^{\text{tr}}$. In contrast, WeightCaster seeks to fit one model $\theta_{t}$ for each ring, such that $\hat{y}=f_{\theta_{t}}(x),\forall x\in X_{t}^{\text{tr}}$. We achieve this via weight space sequence modelling.
2.2 Weight Space Sequence Modelling
The domain decomposition above naturally provides a sequential ordering for the weights $\theta_{t}$, which we use to formulate an initial value problem (IVP). Specifically, our optimisation procedure is the weight space learning problem defined as follows [^2] Note the use of the sum and not the average; this is because some rings might be empty.
@@ 2.3 Stochastic Framework for Regression
To handle uncertainty, we extend the weight space sequence model to a stochastic framework. Instead of a point estimate $\theta_{t}$, the $G_{\phi}$ outputs the parameters of a distribution over weights. This approach allows us to propagate uncertainty from the weight space to the prediction space.
@@ 3.1 Experimental Setup
A classic 1D regression problem where $y=\cos(10x)+0.5x+\epsilon$, where $\epsilon\sim\mathcal{N}(0,25e-6)$. The data is partitioned into disjoint train and test sets. This task requires the model to extrapolate the trend and periodicity into the unseen intervals. We set $\beta=1e-2$. The anchor $\underline{x}$ is chosen as the average across $X^{\text{tr}}$. We specify $T=600$ rings to simulate the weight space sequence, of which $T_{\text{tr}}=300$ overlap the training data.
@@ 4 Conclusion
In this paper, we introduced WeightCaster , a framework that transforms out-of-support generalisation into a weight-space sequence forecasting task. By decomposing the input space into concentric rings and modelling the trajectory of optimal weights, we enable reliable extrapolation with no inductive biases. Our stochastic formulation further provides uncertainty estimates through model linearisation and output-space KL regularization. Experimental results on periodic synthetic data and real-world air quality sensors confirm that WeightCaster outperforms standard regression baselines and competitive generalisation methods. Future work will explore the theoretical underpinnings of this approach in the infinite-length $T\rightarrow\infty$ regime, along with the scaling to high-dimensional manifold data. We will also explore proven strategies for more confident in-distribution uncertainties while simultaneously requiring conservative out-of-support estimates.
@@
By enabling generalisation to unseen scenarios, this work could contribute to the mitigation of catastrophic failures in critical sectors such as environmental monitoring, healthcare, and infrastructure management. Our approach encourages a more transparent understanding of a model’s operational limits, fostering the responsible deployment of predictive AI. To accelerate usability and foster reproducibility, we provide our code at https://anonymous.4open.science/r/oosseq.

### v3

# Out-of-Support Generalisation via Weight-Space Sequence Modelling

@@ Abstract
As breakthroughs in deep learning transform key industries, models are increasingly required to extrapolate on datapoints found outside the range of the training set, a challenge we coin as out-of-support (OoS) generalisation. However, neural networks frequently exhibit catastrophic failure on OoS samples, yielding unrealistic but overconfident predictions. We address this challenge by reformulating the OoS generalisation problem as a sequence modelling task in the weight space, wherein the training set is partitioned into concentric shells corresponding to discrete sequential steps. Our WeightCaster framework yields plausible, interpretable, and uncertainty-aware predictions without necessitating explicit inductive biases, all the while maintaining high computational efficiency. Emprical validation on a synthetic cosine dataset and real-world air quality sensor readings demonstrates performance competitive or superior to the state of the art. By enhancing reliability beyond in-distribution scenarios, these results hold significant implications for the wider adoption of artificial intelligence in safety-critical applications.
@@ 1 Introduction
Leveraging recent advances in sequence modelling and weight-space learning Schürholt et al. [2024]; Nzoyem et al. [2026], we propose WeightCaster : a framework recasting OoS generalisation tasks as forecasting problems through the partitioning of the training set into concentric shells that we call “rings”. By mapping each ring to a discrete time step, we learn sequential dynamics extrapolatable to the test set. Our contributions are threefold: (1) a computationally efficient, parametric, interpretable, and inductive bias-free framework for OoS generalisation; (2) a linearisation strategy enabling the framework to provide uncertainty estimates both in-distribution (InD) and OoS; and (3) empirical validation on synthetic sinusoidal and real-world air quality experiments revealing superior or competitive performance at low parameter count.
@@ 1.2 Related Work
Our approach is also heavily connected to the nascent area of weight-space learning (WSL) Schürholt et al. [2024]. Remarking on the growing capacity of public model repositories such as CivitAI and HuggingFace Jain [2022], WSL has recently captured the attention of the deep learning community, showcasing huge breakthroughs in implicit neural representations Dupont et al. [2022], generalisation error prediction Unterthiner et al. [2020], and sequence modelling Nzoyem et al. [2026].
@@ 2 Method
Fig. 1 summarises our proposed WeightCaster using a 1D sine wave as our example. We begin by selecting an anchor point from the training dataset, and we decompose the training domain into successive hyperspherical shells (equivalent to intervals in 1D or annuli in 2D). While their geometric structure might be different depending on $D_{x}$ and the chosen distance metric, we refer to these shells as “rings” for simplicity. Each ring corresponds to a step in a sequence model, and a weight-space sequence model learns to predict suitable weights for each step.
Figure 1: Illustration of the two main steps of the WeightCaster framework for sinusoidal extrapolation. (a) First, an anchor point is chosen and the input domain $\mathbb{R}^{1}$ is decomposed into $T=10$ “rings”, here clearly delineated as intervals. Within each ring, we consider a simple linear model $\hat{y}=\theta^{1}\cdot x+\theta^{2}$, where $\theta=[\theta^{1},\theta^{2}]^{T}$ contains the slope and intercept to anchor, respectively. (b) Optimal weights $\{\theta_{t}\}_{t=1}^{T_{\text{tr}}}$ for the data in each ring are subsequently computed by fitting a weight-space sequence model, one ring corresponding to one time step. Suitable weights for OoS datapoints are obtained by rolling out the sequence model for time steps $t\geq T_{\text{tr}}$ (in this example, $T_{\text{tr}}=9$).
In the following paragraphs of this section, we detail the ingredients of the WeightCaster framework. We begin by outlining the input domain decomposition and weight-space sequence modelling strategies. Deterministic training and inference algorithms are discussed, followed by a stochastic framework for uncertainty estimation.
@@ 2.1 Domain Decomposition
Conventional machine learning seeks to learn a single model $\hat{y}=f_{\theta}(x)$, $\forall x\in X^{\text{tr}}\triangleq\bigcup_{t=1}^{T_{\text{tr}}}X_{t}^{\text{tr}}$. In contrast, WeightCaster seeks to fit one model $\theta_{t}$ for each ring, such that $\hat{y}=f_{\theta_{t}}(x),\forall x\in X_{t}^{\text{tr}}$. We achieve this via weight-space sequence modelling.
2.2 Weight-Space Sequence Modelling
The domain decomposition above naturally provides a sequential ordering for the weights $\theta_{t}$, which we use to formulate an initial value problem (IVP). Specifically, our optimisation procedure is the weight-space learning problem defined as follows [^2] Note the use of the sum and not the average; this is because some rings might be empty.
@@ 2.3 Stochastic Framework for Regression
To handle uncertainty, we extend the weight-space sequence model to a stochastic framework. Instead of a point estimate $\theta_{t}$, the $G_{\phi}$ outputs the parameters of a distribution over weights. This approach allows us to propagate uncertainty from the weight space to the prediction space.
@@ 3.1 Experimental Setup
A classic 1D regression problem where $y=\cos(10x)+0.5x+\epsilon$, where $\epsilon\sim\mathcal{N}(0,25e-6)$. The data is partitioned into disjoint train and test sets. This task requires the model to extrapolate the trend and periodicity into the unseen intervals. We set $\beta=1e-2$. The anchor $\underline{x}$ is chosen as the average across $X^{\text{tr}}$. We specify $T=600$ rings to simulate the weight-space sequence, of which $T_{\text{tr}}=300$ overlap the training data.
@@ 4 Conclusion
In this paper, we introduced WeightCaster , a framework that transforms out-of-support generalisation into a weight-space sequence forecasting task. By decomposing the input space into concentric rings and modelling the trajectory of optimal weights, we enable reliable extrapolation with no inductive biases. Our stochastic formulation further provides uncertainty estimates through model linearisation and output-space KL regularization. Experimental results on periodic synthetic data and real-world air quality sensors confirm that WeightCaster outperforms standard regression baselines and competitive generalisation methods. Future work will explore the theoretical underpinnings of this approach in the infinite-length $T\rightarrow\infty$ regime, along with the scaling to high-dimensional manifold data. Chosing an appropriate anchor location $\underline{x}$ remains a challenge we wish to solve. We will also explore proven strategies for more confident in-distribution uncertainties while simultaneously requiring conservative out-of-support estimates.
@@
By enabling generalisation to unseen scenarios, this work could contribute to the mitigation of catastrophic failures in critical sectors such as environmental monitoring, healthcare, and infrastructure management. Our approach encourages a more transparent understanding of a model’s operational limits, fostering the responsible deployment of predictive AI. To accelerate usability and foster reproducibility, we provide our code at https://github.com/ddrous/weightcaster.